### v1

@@ 5.3 Synthetic nonlinear MPEC benchmark
Many problems in MacMPEC are relatively small or easy to solve with any of the examined algorithms. Moreover, two-thirds of problems have an empty biactive and very few non-S stationary points, making the verification of B-stationarity often trivial. To compare MPECopt further, we consider a synthetic nonlinear MPEC benchmark, where the solutions have many biactive constraints. Details on the problem generation are described in the Appendix A, and an open-source implementation can be found at https://github.com/nosnoc/mpecopt/benchmarks/nonlinear_mpec.
For the sake of brevity, we only do here the same experiments as in Sections 5.2.3 and 5.2.4. All others can be recreated by running the code in https://github.com/nosnoc/mpecopt/benchmarks/nonlinear_mpec. Notably, in contrast to MacMPEC, in the case of Phase I algorithms, Reg-Simple is slower and less successful than Reg-LPEC on this test set. Moreover, the feasibility $\ell_{1}$ and $\ell_{\infty}$ become more competitive and are closer to Reg-Simple. Overall, Reg-LPEC and $\ell_{1}$-LPEC are still the fastest and most successful Phase I strategies.
Figure 14: Number of NLPs (top plot) and NLPs (bottom plots) solved in the synthetic test set on all problem instances.
Fig. 13 shows the relative and absolute performance profiles, and Fig. 14 the total number of NLPs and LPECs solved on each problem instance. Simile to the results on MacMPEC (Fig. 10), we see that the MPECopt variants are the fastest and most successful algorithms, for the same reasons as in the previous experiments, i.e., a smaller number of total NLP solves is needed, see Fig. 14. Moreover, on this set, the differences in terms of computation times, between MPECOpt-Reg-Gurobi (96.16% success rate), MPECOpt-$\ell_{1}$-Gurobi (96.67% success rate), and Pen-$\ell_{1}$ (96.67% success rate) are less significant, wheres Reg (96.67% success rate) needs to solve more NLPs and is thus solver. MPECOpt-Reg-HiGHS (96.67% success rate) is slightly slower than MPECopt variants using Gurobi, but is still faster than Reg. In cases where they fail, all algorithms report to find a locally infeasible point. It can be seen in Fig. 14 that MPECopt variants solve fewer NLPs than Reg, whereas MPECOpt-$\ell_{1}$-Gurobi and Pen-$\ell_{1}$ usually require the same number of NLP solves. In most cases only one to two LPEC solves are needed, which do not add significantly to the overall computation time, cf. Fig. 13a.
Figure 15: Distribution of stationary points on the synthetic test set for different solution methods. Failure counts the number of infeasible problems.
Next, we examine the different types of stationary points in Fig. 15. This problem set is more balanced, as only about a quarter of the problems have no biactive constraints. To determine the type of stationary point or to verify B-stationarity we use the same iterative procedure as in Section 5.2.4. Since there are many biactive constraints, and MPEC-LICQ is often violated, there are more non-S stationary points. In this set, Pen-$\ell_{1}$ manages to find in all cases B-stationary point, even though MPEC-LICQ is violated. To the best of our knowledge, currently existing theoretical results do not fully explain this. In four problems Reg finds W-stationary points that are not B-stationary. On the other hand, its MPECopt counterpart finds B-stationary points with a significantly better objective value in these cases. We remind the reader that verifying the B-stationarity of non-S-stationary points computed by Reg and Pen-$\ell_{1}$ usually requires solving several additional NLPs and LPECs in addition to the computations shown in Figures 13 and 14. Our experiments empirically confirm, that Reg sometimes does not converge to B-stationary points. However, this did not happen very often in all of our experiments.
@@ 6 Conclusions and summary
This paper introduced MPECopt, a numerical method for computing B-stationary points of mathematical programs with equilibrium constraints (MPECs). MPECopt operates in two phases. The first phase identifies a feasible point of the MPEC or certifies local infeasibility. The second phase iterates over feasible points to find a B-stationary point by solving a finite sequence of branch nonlinear programs (BNLPs) and linear programs with equilibrium constraints (LPECs). The LPEC is directly equivalent to the definition of B-stationarity.
We proposed two distinct strategies for Phase I. The first strategy leverages the effectiveness of existing regularization and penalty-based methods. The second approach involves solving a feasibility MPEC directly using the Phase II algorithm. The paper includes a detailed theoretical analysis of two LPEC subproblem formulations and provides a comprehensive convergence theory for MPECopt. In particular, Theorem  4.2 proves that the algorithm terminates after a finite number of iterations. All theoretical results are illustrated and motivated by a series of tutorial examples.
We argue that the nonconvex LPEC subproblems can be efficiently solved using mixed-integer linear programming (MILP) methods. On two moderate-size benchmark problem sets we demonstrate that this is indeed the case and that solving LPECs is not a computational bottleneck in practice. Interestingly, we prove that if a point is S-stationary, solving a single linear program (LP) is sufficient for solving the MILP reformulation of the LPEC constructed at this point. Another key finding, proven in Theorem 4.1, is that, given a point close enough to a feasible point of the MPEC, then just by finding a feasible point of an LPEC constructed at this point is sufficient for identifying a feasible BNLP. Moreover, the trust-region radius of the LPECs in Phase II can be arbitrarily small, leading to very tight relaxations and efficient solutions via MILP methods – often two order magnitudes faster than homotopy-based methods for LPECs.
@@
In addition to the theoretical analysis, this paper is accompanied by an open-source implementation of MPECopt available at https://https://github.com/nosnoc/mpecopt, including the algorithmic variations discussed. The method is relatively easy to implement, benefiting from robust and mature NLP and MILP solvers. We tested MPECopt on two extensive numerical benchmarks: the MacMPEC set and a synthetic nonlinear MPEC test set, introduced in this paper. Numerical results confirm the efficacy of the proposed methods for finding B-stationary points. MPECopt is usually faster and more reliable than existing methods, requiring fewer NLP solves than existing homotopy-based methods.
@@
An intriguing observation from the benchmarks is that regularization and penalty-based methods frequently identify B-stationary points that are not S-stationary. To our knowledge, the current convergence results for these methods do not explain this phenomenon. However, verifying this currently requires solving additional nonlinear programs and LPECs, whereas MPECopt inherently certifies B-stationarity as solving LPECs is integrated into the algorithm.
Finally, in this paper we assume that the BNLPs are solved to full convergence. A natural direction for future work would be to study inexact BNLP solutions while maintaining robust and fast convergence. This extension would likely require a more sophisticated globalization strategy, since most iterates would not be exactly feasible. However, Theorem 4.1 suggests that LPECs constructed near feasible points have similar properties to those constructed at feasible points. This insight should be exploited in the design of new numerical methods. From an implementation point of view, it would be interesting to test tailored LPEC methods Fang2010 ; Hu2012 ; JaraMoroni2020 ; JaraMoroni2018 within MPECopt.
@@ Acknowledgments
A.N. acknowledges the generous support of this research by the DFG via Research Unit FOR 2401, project 424107692, 504452366 (SPP 2364), and 525018088, by BMWK via 03EI4057A and 03EN3054B, and by the EU via ELO-X 953348. Moreover, A.N. thanks Anton Pozharskiy for translating MacMPEC to CasADi, and for many useful remarks on the MPECopt and benchmark implementations.
Appendix A Generating a synthetic MPEC benchmark
@@
To generate a nonlinear MPEC, we make the following modification to the LPEC (39). We replace the objective in (39a) by $f(x)=f^{\mathrm{nl}}(x_{0},x_{1})+\alpha g_{0}^{\top}x_{0}+\alpha g_{1}^{\top}x_{1}$, where $f^{\mathrm{nl}}(x_{0},x_{1})$ is a nonlinear objective test function from Andrei2008 . Let $c^{\mathrm{ineq}}(x)=Ax_{0}+Bx_{1}-a\geq 0$ define the inequality constraints. To introduce more nonlinearity, we consider a subset of inequality constraint indices, i.e., ${\mathcal{I}}^{\mathrm{ineq}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$. Next, all inequality constraints $c_{i}^{\mathrm{ineq}}(x)\geq 0$ with $i\in{\mathcal{I}}^{\mathrm{ineq}}$ are replaced by $c_{i}^{\mathrm{ineq}}(x)+c_{i}^{\mathrm{ineq}}(x)^{2}+c_{i}^{\mathrm{ineq}}(x)^{4}\geq 0$. Moreover, we duplicated a subset of the inequality constraints (both linear and nonlinear), with indices $i\in{\mathcal{I}}^{\mathrm{d}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$. This makes the MPEC more degenerate as it increases the probability of violating MPEC-LICQ if inequality constraints are active. Next, we define equality constraints by introducing a slack variable $x_{2}\in{\mathbb{R}}^{m}$ for the left hand side in  (39c), which reads as $c^{\mathrm{eq}}(x)=x_{2}-Mx_{0}+Nx_{1}+q$. Analogously to the inequality constraints, we replace some constraints by $c^{\mathrm{eq}}_{i}(x)=0$ by $c^{\mathrm{eq}}_{i}(x)+c^{\mathrm{eq}}_{i}(x)^{2}+c^{\mathrm{eq}}_{i}(x)^{4}=0,i\in{\mathcal{I}}^{\mathrm{eq}}\subseteq\{1,\ldots,m\}$. Moreover, we add upper bounds $x^{\mathrm{ub}}_{0},x^{\mathrm{ub}}_{1}$ on $x_{0}$ and $x_{1}$.
The full pseudocode for generating a single problem instance is given in Algorithm 4. Thereby, $\mathcal{N}(\mu,\sigma^{2})$ denotes the normal distribution with $\mu$ mean and $\sigma$ standard deviation. The elements of $\mathcal{U}_{n}(a,b)$ are $n$ dimensional vectors whose elements are drawn from a uniform distribution on the interval $[a,b]$. Similarly, the elements of $\mathcal{U}_{n\times k}(a,b,s)$ are $n\times k$ matrices whose elements take values from a uniform distribution on $[a,b]$. The scalar $s\in\left(0,1\right]$ is the density of this random matrix, which is defined as the ratio between the number of nonzero elements and the total number of elements of this matrix.
@@
5:Generate upper bounds $x_{0}^{\mathrm{ub}}\sim\mathcal{U}_{p}(100,1000)$ and $x_{1}^{\mathrm{ub}}\sim\mathcal{U}_{m}(100,1000)$.
6:Equality constraints.
7:Choose density for the random matrices $s\sim\mathcal{U}_{1}(0.1,0.3)$.
@@
13:Replace $c_{i}^{\mathrm{eq}}(x)=0$ by $c_{i}^{\mathrm{eq}}(x)+c_{i}^{\mathrm{eq}}(x)^{2}+c_{i}^{\mathrm{eq}}(x)^{4}=0$ for all $i\in{\mathcal{I}}^{\mathrm{eq}}$.
14:Inequality constraints
@@
20:Pick a random subset ${\mathcal{I}}^{\mathrm{d}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$ with cardinality $|{\mathcal{I}}^{\mathrm{d}}|=\mathrm{round}(s^{\mathrm{d}}n_{\mathrm{ineq}})$ and duplicate inequality constraints $c_{i}^{\mathrm{ineq}}(x)\geq 0$ for all $i\in{\mathcal{I}}^{\mathrm{d}}$.
21:Objective function.

### v2

@@ 5.3 Synthetic nonlinear MPEC benchmark
Many MacMPEC problems are small or easily solved by all examined algorithms. To further compare MPECopt, we use a synthetic nonlinear MPEC benchmark described in Appendix D. The benchmark consists of 75 significantly larger problems, ranging from 200 complementarity constraints and 500 variables in the smallest instances to 4000 complementarity constraints and 10000 variables in the largest. These problems are specifically designed to violate the MPEC-LICQ. In this experiment, we exclude the MINLP from the comparison, as it was already not competitive on the easier problem set. Fig. 7 shows the relative and absolute performance profiles of the methods we compared.
(a) Solving to optimality.
(b) Early termination.
Figure 8: Number of nodes in the MILP solver in Phase I and $I$. The node counter on the horizontal axis is on logarithmic scale.
Similar to the results on MacMPEC (Fig. 5), we see that the MPECopt variants are the fastest and most successful algorithms even on the large problem instances. For the same reasons as in the previous experiments, they require fewer total NLP solves. Moreover, on this set, the differences in computation times between MPECOpt-Reg-Gurobi (96.16% success rate), MPECOpt-$\ell_{\infty}$-Gurobi (94.67% success rate), and Pen-$\ell_{\infty}$ (94.67% success rate) are more pronounced, whereas Reg (94.67% success rate) needs to solve more NLPs and is therefore slower. The lower success rates of the regularization- and penalty-based methods compared to their MPECopt counterparts are due to either reaching the maximum number of iterations or hitting the time limit. In other cases, the methods failed by converging to a point of local infeasibility, i.e., when the restoration phase failed. Directly solving the NLP is by far the least successful approach (16.00% success rate), and in most cases it timed out after reaching the 45-minute limit. This further justifies the development of tailored methods for solving MPECs. As shown in Fig. 7a, the other methods compute solutions considerably faster and rarely run into timeouts.
Lastly, we examine the effects of early termination in the LPEC solvers in Fig. 8. In Fig. 8a, where the MILP (and thus the LPEC) is solved to global optimality, we see that in both phases there are instances requiring a significant number of LPs. By contrast, Fig. 8b shows the effects of early termination in Phase I (due to Corollary 4.1) and Phase II (due to Corollary 4.2), where the number of LPs is drastically reduced. In particular, in Phase I a single LP was always sufficient. In Phase II, we also observe a reduction. We terminate the MILP solver if a feasible point with an objective below $0.1\|\nabla f(x^{k})\|\rho^{k,l}$ is found. With more careful tuning of this factor, the number of LPs in Phase II can often be reduced further.
In conclusion, we see that the proposed method is competitive with other existing approaches. Moreover, the experiments demonstrate that solving an LPEC was never a computational bottleneck, even for the larger problem instances. Significant savings can be achieved through early termination, in line with our theoretical predictions.
@@ 6 Conclusions and summary
This paper introduces MPECopt, a two-phase method for computing B-stationary points of mathematical programs with complementarity constraints (MPECs). The first phase identifies a feasible point of the MPEC or certifies local infeasibility. The second phase iterates over feasible points to find a B-stationary point by solving a finite sequence of branch nonlinear programs (BNLPs) and linear programs with complementarity constraints (LPECs). We provide a detailed theoretical analysis of the two LPEC subproblems and prove that the algorithm terminates in a finite number of iterations.
@@
We also provide an open-source implementation of MPECopt, available at github.com/nosnoc/mpecopt. Our implementation leverages robust and mature NLP and MILP solvers. We tested MPECopt on two extensive numerical benchmarks: the MacMPEC set and a new synthetic nonlinear MPEC test set, containing problems with up to four thousand complementarity constraints. The numerical results confirm the efficacy of the proposed method for finding B-stationary points. MPECopt is generally faster and more reliable than existing methods, requiring fewer NLP solves than homotopy-based approaches.
In general, LPECs are challenging combinatorial optimization problems that can be formulated as mixed-integer linear programs (MILPs). Interestingly, a combinatorial number of LPs only needs to be solved if the limit point is B-stationary but not S-stationary. We prove that in all other cases, the LPEC solver can be terminated early or that solving a single LP is sufficient. In our numerical experiments on the two benchmarks, even for large-scale problem instances, it was often sufficient to solve only the root node of the MILP reformulation of the LPEC, which corresponds to a single LP solve.
@@
An intriguing observation from the benchmarks is that the limit point of regularization- or penalty-based methods is in most cases B-stationary, even when it is not S-stationary. To our knowledge, current convergence results for these methods do not explain this phenomenon. However, verifying B-stationarity typically requires solving additional nonlinear programs and LPECs, whereas MPECopt inherently certifies it, since solving LPECs is fully integrated into the algorithm.
Finally, in this paper we assume that the BNLPs are solved to full convergence. A natural direction for future work is to study inexact BNLP solutions while maintaining robust and fast convergence.
@@ Acknowledgments
A.N. acknowledges the generous support of this research by the DFG via Research Unit FOR 2401, project 424107692, 504452366 (SPP 2364), and 525018088, by BMWK via 03EI4057A and 03EN3054B, and by the EU via ELO-X 953348. Moreover, A.N. thanks Anton Pozharskiy for translating MacMPEC to CasADi, and for many useful remarks on the MPECopt and benchmark implementations. This work was also supported by the Applied Mathematics activity within the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under contract number DE-AC02-06CH11357.
Appendix A Proofs of lemmata from Section 3
Proof of Lemma 3.1: Denote by $\Omega^{\mathrm{R}}$ and $\Omega^{\mathrm{F}}$ the feasible sets of the reduced LPEC (12) and full LPEC (15), excluding the constraints $\|d\|\leq\rho$, respectively. Compared to $\Omega^{\mathrm{R}}$, the set $\Omega^{\mathrm{F}}$ is additionally restricted by the linear inequalities for all inactive constraints, i.e., $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,i\in\bar{{\mathcal{A}}}^{k}$. Recall that the constraints $0\leq x_{1,i}^{k}+d_{1,i}\perp x_{1,i}^{k}+d_{2,i}\geq 0$ define the set $\{d\mid x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0\}\cup\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}$. Moreover, $\Omega^{\mathrm{F}}$ contains all branches excluded for the strictly active complementarity indices $i\in{\mathcal{I}}_{+0}^{k}\cup{\mathcal{I}}_{0+}^{k}$ in $\Omega^{\mathrm{R}}$. Therefore, the two sets are related as follows:
$\begin{split}\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}&\bigcup_{i\in{\mathcal{I}}_{0+}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0\}\\ &\bigcup_{i\in{\mathcal{I}}_{+0}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}\\ &\bigcap_{i\in\bar{{\mathcal{A}}}^{k}}\{d\mid c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0\}.\end{split}$ (39)
To ensure that the two LPECs have the same set of solutions, we need to restrict $\|d\|_{\infty}\leq\rho$ such that $\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}$.
Let us first consider the inequality constraints. For all $i\in\bar{{\mathcal{A}}}^{k}$, we have $c_{i}(x^{k})>0$, and we want $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d>0.$ If $\nabla c_{i}(x^{k})=0$, this inequality holds trivially for any $d$ (and thus any $\rho$). Otherwise, if $\nabla c_{i}(x^{k})\neq 0$, we can apply the Cauchy–Schwarz inequality $|\nabla c_{i}(x^{k})^{\top}d|\leq\|\nabla c_{i}(x^{k})\|\,\|d\|.$ Hence, the linearized inequality remains inactive as long as $\|d\|<\frac{c_{i}(x^{k})}{\|\nabla c_{i}(x^{k})\|}.$ To ensure this holds for all relevant $i$, we define $\bar{\rho}_{1}=\min_{i\in\bar{{\mathcal{A}}}^{k},\,\nabla c_{i}(x^{k})\neq 0}\frac{c_{i}(x^{k})}{\|\nabla c_{i}(x^{k})\|}.$ Next, we compute upper bounds for $\rho$ to exclude the union terms in (39).
Now observe that, for example, for any $i\in{\mathcal{I}}_{+0}^{k}$ it holds that $x_{1,i}^{k}>0$ and $x_{2,i}=0$. If $\rho$ is chosen such that $d_{1,i}<\rho<x_{1,i}^{k}$ then $x_{1,i}^{k}+d_{1,i}>0$ and $0\leq x_{1,i}^{k}+d_{1,i}\perp d_{2,i}\geq 0$ can be only satisfied for $d_{2,{i}}=0$. Thus, the branch $\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}$ is infeasible. Similar reasoning holds for any $i\in{\mathcal{I}}_{0,+}^{k}$. Therefore, by picking $\rho<\bar{\rho}_{2}=\min\{\{x_{1,i}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{x_{2,i}\mid i\in{\mathcal{I}}_{0+}^{k}\}\}$, none of the terms in the unions in (39) is feasible. The equations (16) is follows from writing $\bar{\rho}=\min(\bar{\rho}_{1},\bar{\rho}_{2})$ more explicitly. Thus, for $\rho<\bar{\rho}$, we have $\Omega^{\mathrm{F}}\cap\{\|d\|_{\infty}\leq\rho\}=\Omega^{R}\cap\{\|d\|_{\infty}\leq\rho\}$. As the two LPECs have the same objective, their sets of local minimizers coincide. ∎
Proof of Lemma 3.2: It follows from Lemma 3.1 that for $\rho\in(0,\bar{\rho})$ both the full and reduced LPEC have the same feasible set.
Thus, we focus solely on the reduced LPEC. Recall from Eq. (10) that $\Omega=\bigcup_{(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})\in\mathcal{P}^{k}}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})}}$, where $\mathcal{P}^{k}$ denotes the set of all possible partitions $(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})$ at $x^{k}\in\Omega$, cf. (7). This relation also holds for the tangent cones of the MPEC and the BNLPs, as well as for the MPEC-linearized feasible cone and the standard linearized feasible cones of the BNLPs (Flegel2005a, , Lemma 3.2). The latter reads explicitly as:
$\begin{split}\mathcal{F}_{\Omega}(x^{k})=\bigcup_{(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})\in\mathcal{P}^{k}}\{d\in{\mathbb{R}}^{n}\mid&c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,\ \forall i\in{\mathcal{A}}^{k},\\ &x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0,\ \forall i\in{\mathcal{I}}_{1}^{k},\\ &x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0,\ \forall i\in{\mathcal{I}}_{2}^{k}\}.\end{split}$ (40)
[…]
Appendix B Identifying stationary points
To determine the type of stationary point from Definition 2.3 we must solve the TNLP (9), which requires a correct identification of the active sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+}$ and ${\mathcal{I}}_{00}$. MPECopt is an active set method where the complementarity constraints are usually satisfied exactly or close to machine precision, making it easy to find the active sets for the TNLP.
On the other hand, Reg, and Pen-$\ell_{1}$ rarely have a complementarity residual $h_{\perp}(x)$ much smaller than the prescribed tolerance of $10^{-9}$, making an active set identification more difficult. We take an iterative procedure, and identify the active set via:
${\mathcal{I}}_{00}^{\epsilon}(x)=\{i\mid x_{1,i}\leq\epsilon,x_{2,i}\leq\epsilon\},$ (41c)
with $\epsilon=10^{-6}$. To minimize the possibility of wrong conclusions, we implement several safeguards. We denote the solution computed via Reg or Pen-$\ell_{1}$ by $x^{*}$, and the corresponding TNLP solution by $x^{\mathrm{tnlp}}$. We declare the active set to be correctly identified if $\|x^{*}-x^{\mathrm{tnlp}}\|\leq 10^{-7}$. If this is not satisfied we reduce $\epsilon$ by a factor of $10$ and repeat the procedure. Once $\|x^{*}-x^{\mathrm{tnlp}}\|\leq 10^{-7}$ is true or a maximum number of 5 attempts is exceeded, we compute the multipliers in the TNLP and classify the stationary point. To verify B-stationarity, we take several steps. First, we use the point $x^{\mathrm{tnlp}}$ and construct an ${\mathrm{LPEC}}(x^{\mathrm{tnlp}},\rho)$ with $\rho=10^{-3}$. If the objective of this LPEC is below $\texttt{tol}_{\mathrm{B}}=10^{-8}$, then we declare a $x^{*}$ to be B-stationary. If this is not the case, we take up to three more LPEC solve attempts while reducing $\rho$ by a factor of 10. The minimal value is $\rho=10^{-6}$, which is the minimal allowed trust region radius in MPECopt. This delivers a valid certificate in most cases. To have more robustness to wrong active set identification, if this does not verify B-stationarity, we repeat the same but now with $x^{*}$ as linearization point ${\mathrm{LPEC}}(x^{*},\rho)$. Finally, if this also fails, we look at the objective $f(x^{*})$ and the objective computed by MPECopt for the same problem, which is denoted by f(x^)f($\hat${x}). If the relative difference satisfies $\frac{|f(x^{*})-f(\hat{x})|}{|f(\hat{x})|+10^{-16}}<10^{-3}$, then we declare $x^{*}$ to be B-stationary, if $\hat{x}$ is B-stationary. If the difference between the objective is larger, and the LPECs find a descent direction, the point $x^{*}$ is declared to be not B-stationary.
Appendix C Comparison of different LPEC solvers
We compare the cumulative NLP and LPEC computation times in every MPECopt solver call on every MacMPEC problem. In all experiments, we use the regularization-based Phase I described in Algorithm 2, but with different LPEC solvers in both phases. We compare solving the LPEC as MILP (18) and via regularization-based methods. The HiGHS solver performs significantly better if the trust region in the LPEC is smaller, thus we set $\rho_{I}=10^{-3}$ in Phase I, while Gurobi is less sensitive to $\rho_{I}$ values. We consider four variants: Gurobi-MILP ($\rho_{I}=10^{-1}$), HiGHS-MILP ($\rho_{I}=10^{-3}$), Reg-MPEC, and $\ell_{\infty}$-MPEC.
We solve all MacMPEC problems with four different LPEC solution strategies, comparing total cumulative computation times for solving all LPECs in each instance. In 34 of 191 problems, the MPEC is already solved or locally infeasible after the first NLP, so we exclude these (effective LPEC time is zero). The remaining 157 problems are solved by all approaches except HiGHS-MILP, which fails on siouxfls and siouxfls1 due to unsolved LPECs.
Fig. 9 shows the results, where the diagonal represents equal NLP and LPEC solving times. For MILP methods, most points fall below the diagonal, indicating that BNLPs dominate computation time rather than LPECs. For regularization methods, points cluster around/above the diagonal, showing computational burden is split almost equally between NLPs and LPECs. Specifically, cumulative NLP time exceeds LPEC time in 0% of cases for Gurobi-MILP, 20% for HiGHS-MILP, 87.26% for Reg-MPEC, and 85.99% for $\ell_{\infty}$-MPEC. MILP solvers benefit from smaller initial trust region radii in Phase I, but smaller $\rho_{I}$ requires solving more Phase I NLPs to find feasible LPECs, shifting computational load toward NLPs.
Somewhat surprisingly, the global MILP algorithms significantly outperform the local regularization and penalty-based methods. The MILP-based methods are up to two orders of magnitude faster than regularization-based methods for LPECs on the MacMPEC test set. This difference can be explained by several observations. The MILP methods solve the LPEC to global optimality (sufficient for verifying B-stationarity), whereas the others compute only a stationary point (which may not be a local minimum or B-stationary point). Consequently, LPECs solved by regularization-based methods are more often locally infeasible in Phase I or require solving more LPECs until a B-stationary point is verified in Phase II. This confirms the arguments in Section 3.2.2 on efficiently solving LPECs using MILP reformulations. Moreover, when the solution is S-stationary, solving the MILP requires solving a single LP, complying with Proposition 3.1.
This experiment suggests that solving LPECs with MILP methods as in MacMPEC can be done very efficiently in practice, both with commercial (Gurobi) and open-source (HiGHS) solvers.
Appendix D Generating a synthetic MPEC benchmark
@@
To generate a nonlinear MPEC, we make the following modification to the LPEC (42). We replace the objective in (42a) by $f(x)=f^{\mathrm{nl}}(x_{0},x_{1})+\alpha(g_{0}^{\top}x_{0}+g_{1}^{\top}x_{1})$, where $f^{\mathrm{nl}}(x_{0},x_{1})$ is a nonlinear objective test function from unconstrained problems from the CUTEst test set Gould2013 Let $c^{\mathrm{ineq}}(x)=Ax_{0}+Bx_{1}-a\geq 0$ define the inequality constraints. To introduce more nonlinearity, we consider a subset of inequality constraint indices, i.e., ${\mathcal{I}}^{\mathrm{ineq}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$. Next, all inequality constraints $c_{i}^{\mathrm{ineq}}(x)\geq 0$ with $i\in{\mathcal{I}}^{\mathrm{ineq}}$ are replaced by $c_{i}^{\mathrm{ineq}}(x)+c_{i}^{\mathrm{ineq}}(x)^{2}+c_{i}^{\mathrm{ineq}}(x)^{4}\geq 0$. Moreover, we duplicated a subset of the inequality constraints (both linear and nonlinear), with indices $i\in{\mathcal{I}}^{\mathrm{d}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$. This makes the MPEC more degenerate as it increases the probability of violating MPEC-LICQ if inequality constraints are active. Next, we define equality constraints by introducing a slack variable $x_{2}\in{\mathbb{R}}^{m}$ for the left hand side in  (42c), which reads as $c^{\mathrm{eq}}(x)=x_{2}-Mx_{0}+Nx_{1}+q$. Analogously to the inequality constraints, we replace some constraints by $c^{\mathrm{eq}}_{i}(x)=0$ by $c^{\mathrm{eq}}_{i}(x)+c^{\mathrm{eq}}_{i}(x)^{2}+c^{\mathrm{eq}}_{i}(x)^{4}=0,i\in{\mathcal{I}}^{\mathrm{eq}}\subseteq\{1,\ldots,m\}$. Moreover, we add upper bounds $x^{\mathrm{ub}}_{0},x^{\mathrm{ub}}_{1}$ on $x_{0}$ and $x_{1}$.
The full pseudocode for generating a single problem instance is given in Algorithm 4. Thereby, $\mathcal{N}(\mu,\sigma^{2})$ denotes the normal distribution with $\mu$ mean and $\sigma$ standard deviation. The elements of $\mathcal{U}_{n}(a,b)$ are $n$ dimensional vectors whose elements are drawn from a uniform distribution on the interval $[a,b]$. Similarly, the elements of $\mathcal{U}_{n\times k}(a,b,s)$ are $n\times k$ matrices whose elements take values from a uniform distribution on $[a,b]$. The scalar $s\in\left(0,1\right]$ is the density of this random matrix, which is defined as the ratio between the number of nonzero elements and the total number of elements of this matrix. We chose the density of the random matrices to be up to 10%. For larger problems with more than 3000 variables, this results in quite dense matrices, and generating the computational graphs for the first and second-order derivatives can take several hours. To ensure that the benchmark can be solved in under a week, we set the density of a random matrix $A\in{\mathbb{R}}^{n\times m}$ to $s=\frac{n}{nm}$ for problems with more than 3000 variables.
@@
5: Generate upper bounds $x_{0}^{\mathrm{ub}}\sim\mathcal{U}_{p}(100,1000)$ and $x_{1}^{\mathrm{ub}}\sim\mathcal{U}_{m}(100,1000)$. 6:Equality constraints.
7: Choose density for the random matrices $s\sim\mathcal{U}_{1}(0.05,0.1)$.
@@
13: Replace $c_{i}^{\mathrm{eq}}(x)=0$ by $c_{i}^{\mathrm{eq}}(x)+c_{i}^{\mathrm{eq}}(x)^{2}+c_{i}^{\mathrm{eq}}(x)^{4}=0$ for all $i\in{\mathcal{I}}^{\mathrm{eq}}$. 14:Inequality constraints
@@
20: Pick a random subset ${\mathcal{I}}^{\mathrm{d}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$ with cardinality $|{\mathcal{I}}^{\mathrm{d}}|=\mathrm{round}(s^{\mathrm{d}}n_{\mathrm{ineq}})$ and duplicate inequality constraints $c_{i}^{\mathrm{ineq}}(x)\geq 0$ for all $i\in{\mathcal{I}}^{\mathrm{d}}$. 21:Objective function.